When the reaction 2H_2S(g)^f 2H_2(g) + S_2(g) is c
Answer #1

let the initial pressure of H2S= P

let x = partial pressure of S2 formed

So at equilibrium, partial pressures : H2S= P-2x, H2= 2x, S2= x

given at equilibrium 2x=0.25, x= 0.125 atm

K=0.012= [S2] [H2]2/ [H2S]2 , =0.125*(0.25)2/ (P-0.25)2 = 0.007813/ (P-0.25)2 =0.012

(P-0.25)2= 0.65

P-0.25 = 0.81

P= 0.81+0.25 =1.06 atm ( A is correct)